OpenStudy (austinl):

Differential Equations Question Find approximate values of the solution of the given initial value problem at t=0.1,0.2,0.3,0.4 using the Euler method with h=0.1 Find the solution \(y=\phi(t)\) of the given problem and evaluate \(\phi(t)\) at t=0.1,0.2,0.3 and 0.4 \(y\prime = 3\cos(t)-2y\)

terenzreignz (terenzreignz):

\[\Large y' = 3\cos(t) - y\]

terenzreignz (terenzreignz):

Well, let's work on the hard part first, actually solving this this differential equation... or have you already done that?

OpenStudy (austinl):

-2y*

terenzreignz (terenzreignz):

Whoopsie \[\Large y' = 3\cos(t) - 2y\] So HAVE you solved this? :D

OpenStudy (austinl):

\(\dfrac{dy}{dt}=3\cos(t)-2y\) \(dy=(3\cos(t)-2y)dt\) \(\int dy=\int (3\cos(t)-2y)dt\) Correct so far?

terenzreignz (terenzreignz):

uhh.. no.... there's a y-term in the right-side... one does not simply integrate that with respect to t...

OpenStudy (austinl):

Well, I am drawing a blank then..... :/

terenzreignz (terenzreignz):

this looks like another one of them differentials that need integrating factors...

terenzreignz (terenzreignz):

\[\Large y' + 2y = 3\cos(t)\]

terenzreignz (terenzreignz):

Integrating factor would be \[\LARGE e^{\int2dt}\]

OpenStudy (austinl):

\(\Large{\mu(t)=e^{2t}}\) \(\Large{e^{2t}y^{\prime}}+2e^{2t}y=e^{2t}3\cos(t)\) Look ok so far?

terenzreignz (terenzreignz):

Yup :)

terenzreignz (terenzreignz):

Left side is \[\Large d(e^{2t}y)= 3e^{2t}\cos(t)dt\]

terenzreignz (terenzreignz):

I'll leave you to do the horrid integration.. :3

OpenStudy (austinl):

Gimme a minute......

terenzreignz (terenzreignz):

<starts stopwatch>

OpenStudy (austinl):

\(\Large{y=\dfrac{3(\cos(t)+\sin(t))}{5}}\)

OpenStudy (austinl):

Alright, so we have solved this lovely little differential equation. What now?

terenzreignz (terenzreignz):

You forgot the constant :3

OpenStudy (austinl):

I am looking at an example, and they aren't showing a constant..... Oh well. \(\Large{y=\dfrac{3(\cos(t)+\sin(t))}{5}+C}\)

terenzreignz (terenzreignz):

Crud... I meant \[\Large y e^{2t}= \frac{3\color{red}{e^{2t}}[\cos(t)+ \sin(t)]}{5}+C\]

terenzreignz (terenzreignz):

Wait a sec... rewind... could you redo that integration? Wolfie doesn't agree with you...

terenzreignz (terenzreignz):

\[\Large d(e^{2t}y)= 3e^{2t}\cos(t)dt\]

OpenStudy (austinl):

I made a typo, but it does agree, \(\color{blue}{2\cos(t)}\)*

terenzreignz (terenzreignz):

Okay, so we have this, aye? \[\Large y e^{2t}= \frac{3\color{}{e^{2t}}[2\cos(t)+ \sin(t)]}{5}+C\]

terenzreignz (terenzreignz):

Now, you had an initial condition?

OpenStudy (austinl):

\(y(0)=0\) Most interesting initial condition ever.

OpenStudy (austinl):

Oh duh, plug that bad boy in and solve or C yeah?

terenzreignz (terenzreignz):

Yup :D

terenzreignz (terenzreignz):

Don't lament y(0) = 0 as an initial condition... it usually simplifies a lot of things XD

OpenStudy (austinl):

\(\Large{ye^{2t}=3e^{2t}\dfrac{[2cos(t)+sin(t)]}{5}-\dfrac{6}{5}}\)

terenzreignz (terenzreignz):

Okay... so more on those later.. I'll let you worry about checking about individual points t = 0.1 and so on... do we proceed to using Euler's method?

OpenStudy (austinl):

Umm..... honestly? I have no clue. I am looking at the example and it make no sense.

terenzreignz (terenzreignz):

What makes no sense? Euler's Method?

OpenStudy (austinl):

Yeah, this text book and how it is written makes no sense. It kinda jumps around a lot.

terenzreignz (terenzreignz):

Okay, well, it's already well set-up for Euler's method... Let's look at the original statement of the problem \[\Large y' = 3\cos(t) - 2y\]

terenzreignz (terenzreignz):

\[\Large y'(t) = f(t,y(t))\] right?

terenzreignz (terenzreignz):

Namely, that the derivative of y with respect to t is a function of t and y. Catch me so far?

OpenStudy (austinl):

I follow me thinks. :)

terenzreignz (terenzreignz):

Okay, first we select a step-size, h. The smaller, the more accurate. The choices of t being 0.1, 0.2, 0.3, 0.4 seem to imply that we're to choose a step-size of 0.1. You okay with that, or you want 0.05 for more accuracy?

terenzreignz (terenzreignz):

Or is accuracy not an issue and/or you are specifically instructed to use a step size of h = 0.1?

OpenStudy (austinl):

Find approximate values of the solution of the given initial value problem at t=0.1,0.2,0.3,0.4 using the Euler method with \(\color{blue}{h=0.1}\) I am guessing that I am supposed to use that.

terenzreignz (terenzreignz):

Okay, well, simpler that way, anyway :)

terenzreignz (terenzreignz):

So, this is the heart and soul of the Euler method: \[\Large y_{n+1} =y_n + hf(t_n,y_n) \]does that look daunting? :D

terenzreignz (terenzreignz):

Since the step size is h = 0.1, then we have \[\Large \left.\begin{matrix}t_0 &0\\t_1&0.1\\t_2&0.2\\t_3&0.3\\t_4&0.4\end{matrix}\right.\]

terenzreignz (terenzreignz):

So, ready to begin?

terenzreignz (terenzreignz):

@austinL don't worry, it's a lot easier than it makes itself out to be... so... do we proceed? :D

OpenStudy (austinl):

Yes, sorry. I went and made some coffee! :D Yep, good to go.

terenzreignz (terenzreignz):

Okay, we want to know the approximation for \(\large y_1\) so, we plug in :) \[\Large y_{n+1} =y_n + hf(t_n,y_n)\] \[\Large y_{1} =y_0 + hf(t_0,y_0)\]Got it?

OpenStudy (austinl):

Yep. :)

terenzreignz (terenzreignz):

This is pretty much the only time where you actually *need* that initial condition. But it's pretty darn important, as without it, you don't have a place to start XD So you know that y(0) = 0 that means \[\Large y_0 =0\]\[\Large t_0=0\] Now solve for \(\large y_1\)

terenzreignz (terenzreignz):

You also might want to have a calculator handy

OpenStudy (austinl):

\(y_1=y_0+hf(t_0,y_0)=0+(0.1)(3cos(0)-2(0))=0.1\) ?

terenzreignz (terenzreignz):

As I recall... 3cos(0) = 3...

OpenStudy (austinl):

Oh duh... silly error. 0.3

OpenStudy (austinl):

2*0=0 Austin, get it together.

terenzreignz (terenzreignz):

Okay, great :D Now you have \(\large y_1\) which is an approximation for \(\large y(t_1)=y(0.1)\) Great:D

terenzreignz (terenzreignz):

Can you predict the next step? Work on \(\large y_2\) \[\Large y_{2} =y_1 + hf(t_1,y_1)\]

terenzreignz (terenzreignz):

Note that it's y1 and t1 now, since you now know the value of y1.

OpenStudy (austinl):

Would it be 0.238501?

terenzreignz (terenzreignz):

I.... hang on :D

terenzreignz (terenzreignz):

How did you do this?

OpenStudy (austinl):

\(t_1=0.1,~~y_1=0.3,~~h=0.1\) Plug it in?

terenzreignz (terenzreignz):

I get 0.658 something

OpenStudy (austinl):

\(\Large{y_2=0.3+(0.1)(3\cos(0.1)-2(0.3))}\) I just got a different answer, but is this the correct format?

terenzreignz (terenzreignz):

Should be.

terenzreignz (terenzreignz):

hang on again...

terenzreignz (terenzreignz):

I definitely get 0.6585 something try it again.

OpenStudy (austinl):

I get 0.538501 What am I doing wrong D:

terenzreignz (terenzreignz):

google https://www.google.com/#q=0.3+%2B+(0.1)(3cos(0.1)+%2B+2(0.3))

OpenStudy (austinl):

Ok, what the deuce...

terenzreignz (terenzreignz):

I really don't know...

terenzreignz (terenzreignz):

that is... why you're not getting it... why don't you list down your steps here, one at a time

OpenStudy (austinl):

For the next term I have, 1.08423

terenzreignz (terenzreignz):

Sorry, what? You used y1 = 0.658something?

terenzreignz (terenzreignz):

And yes, that's right :)

OpenStudy (austinl):

Yep, y2=0.658501

terenzreignz (terenzreignz):

So you finally got it then? :D

OpenStudy (austinl):

I do believe so! Thanks very much!

terenzreignz (terenzreignz):

You have y1, y2, y3 all down. finish it off with y4

terenzreignz (terenzreignz):

@austinL hey don't leave me hanging, buddy D:

OpenStudy (austinl):

I am just finishing up. \(t=~~~~~~~h=0.1\) \(0.0~~~~~~~~~0.3\\ 0.1~~~~~~~~~0.658501\\ 0.2~~~~~~~~~1.08423\\ 0.3~~~~~~~~~1.58768\\ 0.4~~~~~~~~~2.18153\)

terenzreignz (terenzreignz):

a bit off...

terenzreignz (terenzreignz):

2.18153... is actually y5 (or when t = 0.5)

OpenStudy (austinl):

Don't need t=0.5

terenzreignz (terenzreignz):

Earth to austin... when t = 0, y should be also be 0 Like... you know, your initial value? XD

terenzreignz (terenzreignz):

Adjusting... \[0.1~~~~~~~~~0.3\\ 0.2~~~~~~~~~0.658501\\ 0.3~~~~~~~~~1.08423\\ 0.4~~~~~~~~~1.58768\\ 0.4~~~~~~~~~2.18153\]

OpenStudy (austinl):

omfg.... *facepalm*

terenzreignz (terenzreignz):

Good thing I asked for a confirm. But you've pretty much gotten the Euler method down again, right? :D

OpenStudy (austinl):

Yeah, I feel much better about it :)

terenzreignz (terenzreignz):

Whoops... \[0.1~~~~~~~~~0.3\\ 0.2~~~~~~~~~0.658501\\ 0.3~~~~~~~~~1.08423\\ 0.4~~~~~~~~~1.58768\\ \color{red}{0.5}~~~~~~~~~2.18153\\TJ ~~~~~~~~~awsum \ \text{^_^ }\]

OpenStudy (austinl):

Thank you very very very much!!!

terenzreignz (terenzreignz):

No problem... now time for me to catch some sleep :) Nighty night ^_^ ----------------------------- Terence out

OpenStudy (austinl):

Nighty night!